OpenStudy (anonymous):

I GIVE MEDALS Given f(x) = 2x + 3 find (f º f)(x): Select one: a. 2x^2 + 13 b. 4x + 9 c. 4x^2 + 9 d. 4x^2 + 13

OpenStudy (anonymous):

I believe this is just multiplying the equation by itself so it'd be (2x+3)(2x+3)

OpenStudy (anonymous):

so b?

OpenStudy (math&ing001):

(f º f)(x) = f(f(x))

OpenStudy (anonymous):

Looks like c to me

OpenStudy (anonymous):

Because 2x*2x multiplies the x terms and the 2's giving you 4x^2

OpenStudy (math&ing001):

No @JoshR1213 it's not multiplication.

OpenStudy (anonymous):

Oh, is it squaring?

OpenStudy (anonymous):

Still gives you the same answer I gave

OpenStudy (math&ing001):

Just replace x with f(x) like this : (f º f)(x) = f(f(x)) = 2(2x+3) + 3

OpenStudy (anonymous):

so it is c

OpenStudy (anonymous):

cause that would be 4x^2+9

OpenStudy (anonymous):

If you solve it the way he set it up it's b

OpenStudy (math&ing001):

2(2x+3) = 4x + 9

OpenStudy (anonymous):

oh ok so there is no 4x^2 its just 4x

OpenStudy (anonymous):

If he's right, yes

OpenStudy (anonymous):

ok